# The Quasar ULAS J1120+0641 May Never Reach a āNormalā Black Hole to Stellar Mass Ratio

### v1

@@ 1 Introduction
However, the fact remains that such extreme galaxies, with , are not observed in the local Universe aside from a few niche cases where the host galaxy has lost significant mass to tidal stripping. If these quasar host galaxies are to grow enough to land on the observed local relationship between supermassive black hole (SMBH) mass and galaxy mass by , they must be able to form stars and merge with other galaxies in their environment. When we observe quasars at the end of the Epoch of Reionization, we therefore expect to see them living in dense environments, surrounded by other galaxies.
For decades, observations have targeted high-redshift quasars in hopes of detecting galaxy overdensities surrounding them. At high redshift (), HST and ground-based optical/NIR telescopes can only access the rest-frame ultraviolet, and must rely on the Lyman break technique (Adams etĀ al., 2015; Husband etĀ al., 2013; Morselli etĀ al., 2014; Ota etĀ al., 2018; Pudoka etĀ al., 2024; Stiavelli etĀ al., 2005; Zheng etĀ al., 2006) or selection of Lyman-alpha emitters (Chanchaiworawit etĀ al., 2019; Harikane etĀ al., 2019; Hennawi etĀ al., 2015; Kashikawa etĀ al., 2007; Overzier etĀ al., 2009), with relatively few bands longward of the Lyman break, to identify candidate galaxies. Although galaxies were detected around many quasars, it remained difficult to characterize the environments of high-redshift quasars without access to the rest-frame optical at high sensitivity and resolution.
@@ 2.1 NIRCam
We pair each of these NIRCam LW filters with a SW filter: F090W ( = 10522.0 s), F150W ( = 5239.5 s), and F200W ( = 5239.5 s). These data, combined with deep archival HST F814W observations (see Section 2.2) provide additional photometric redshift constraints to exclude lower- and higher-redshift interlopers from our sample. In particular, F090W straddles the Lyman break and F814W lies blueward of it for galaxies near .
@@ 3.1 Candidate Identification
We visually inspected these 523 candidates to narrow the sample before performing SED fitting. 69 of the 523 sources are spurious: these are a mix of detector artifacts, shredded diffraction spikes of bright point sources in the image, and real galaxies only partially overlapping with the image in one or more bands (see the top panels of Figure 1 for an example). We discard all obviously spurious sources, leaving us with 454 candidates.
We performed a final round of visual inspection to remove galaxies clearly at extremely low redshift. 160 of the 454 remaining candidates were discarded based on clearly resolved spiral structure, spatial extent, and/or flux inconsistent with even the most extreme galaxy (1 Jy in the UV continuum bands F090W, F150W, and F200W). For examples of discarded galaxies, see the lower panels of Figure 1.
@@ 3.2 Spectral Energy Distribution Fitting
We ran Prospector under these conditions three times for the 292 candidates identified in Section 3.1, with different tophat redshift priors: (v1), (v2), and finally (v3). Galaxies far from the quasar redshift will be worse-fit under the stricter redshift priors, yielding a larger . On the other hand, galaxies near the quasar redshift are likely to have a smaller value in v3 than in v2 or v1, and/or be fit to in the less restrictive v1 or v2 runs.
@@
We identify 28 sources with a smaller with the most restrictive redshift prior (v3) than in v2 or v1. We also find 5 sources that have higher values in v3, but are also fit to in the lower v1 or v2 runs. Our Prospector-curated sample therefore consists of 35 galaxies total.
@@
For an additional check on the best-fit redshifts and resulting galaxy properties from Prospector, we fit these 35 galaxies with Bagpipes. We ran Bagpipes with a delayed-tau star formation history and Kroupa (2001) initial mass function, as with Prospector. The Kriek & Conroy (2013) dust law is not supported in Bagpipes, so we use the Salim etĀ al. (2018) law, which is parameterized similarly. We allow the maximum age of the stellar population to vary from zero to the age of the Universe at the fit redshift. The total mass of stars formed was allowed to vary from 106 to 1012 Mā, and the metallicity from 0.1% to 100% solar. Modifying the functional form of the star formation history did not significantly affect the resulting stellar mass (5-10% difference) or quality of the fit. For most sources, modifying the dust extinction law to a CF00 or Calzetti law changed the resulting stellar mass by , though in a few cases the mass changed by up to . As with Prospector, we ran Bagpipes three times with three different redshift ranges: (v1), (v2), and finally (v3). The v3 range is larger for Bagpipes than for Prospector, because Bagpipes imposes a hard limit at the edges of the range, while Prospector can fit galaxies to a redshift outside the prior range.
@@
22 of the 35 Prospector candidates are fit to a v3 lower than or comparable to v2 and v1 with Bagpipes. Of the remaining 13, most are better fit to , with the F405N filter on H. Many of these galaxies are brighter in F200W than F150W, corresponding to the wavelength of the Balmer break at . These galaxies also tend to have larger values in the Prospector fits, and tend to lie at the bright end of the Prospector sample. For that reason, we remove these 13 galaxies from the sample and continue with our final sample of 22 galaxies. Their properties are summarized in Table 1.
@@ 3.3 Final Sample
ID | RA (ā) | Dec (ā) | Field | /Mā) | /Mā) | |||
300 | 169.99902 | 6.70370 | Q | 7.11 | 7.10 | 8.36 | 7.43 | |
335 | 169.99734 | 6.69700 | Q | 7.10 | 7.11 | 8.98 | 8.51 | |
337 | 169.99750 | 6.69704 | Q | 7.11 | 7.11 | 9.02 | 7.89 | |
414 | 169.99737 | 6.68958 | Q | 7.10 | 7.12 | 8.71 | 8.15 | |
427 | 169.99444 | 6.68186 | Q | 7.11 | 7.11 | 8.66 | 7.84 | |
534 | 170.00581 | 6.69992 | Q | 7.11 | 7.11 | 9.12 | 8.10 | |
624 | 170.00959 | 6.70138 | Q | 7.11 | 7.11 | 8.56 | 7.71 | |
666 | 170.00589 | 6.68920 | Q | 7.11 | 7.11 | 9.19 | 8.73 | |
703 | 170.00662 | 6.68699 | Q | 7.11 | 7.11 | 9.21 | 7.89 | |
709 | 170.00678 | 6.68705 | Q | 7.07 | 7.10 | 9.26 | 8.43 | |
789 | 170.01207 | 6.69086 | Q | 7.14 | 7.11 | 9.02 | 8.03 | |
876 | 170.01181 | 6.68086 | Q | 7.10 | 7.11 | 8.37 | 7.60 | |
941 | 170.01047 | 6.67159 | Q | 7.13 | 7.11 | 9.00 | 8.05 | |
181 | 169.97607 | 6.66299 | A | 7.11 | 7.08 | 9.12 | 8.08 | |
358 | 169.97943 | 6.65433 | A | 7.11 | 7.08 | 8.96 | 8.41 | |
366 | 169.97875 | 6.65236 | A | 7.11 | 7.08 | 9.34 | 7.96 | |
456 | 169.97744 | 6.64111 | A | 7.12 | 7.08 | 8.83 | 8.16 | |
634 | 169.97744 | 6.62748 | A | 7.09 | 7.08 | 9.48 | 8.25 | |
693 | 169.98564 | 6.64099 | A | 7.11 | 7.08 | 8.89 | 8.07 | |
160* | 169.96941 | 6.64997 | A | 7.23 | 7.31 | 10.41 | 9.26 | |
856* | 169.99940 | 6.65472 | A | 7.01 | 7.41 | 11.25 | 10.12 | |
911* | 169.98969 | 6.62826 | A | 7.20 | 7.17 | 11.29 | 9.60 |
The positions of the 22 candidates are shown, along with the position of the quasar, in Figure 2, and their thumbnails and SEDs are plotted in the Appendix. Thirteen candidates lie in Module B, in the field of the quasar; the remaining nine lie in the adjacent field (Module A). The vast majority of these candidates are fit to a redshift between 7.07 and 7.14 by both Prospector and Bagpipes, but three of the candidates in the adjacent field are fit farther from the quasar redshift by both codes. We consider these three candidates āmedium confidence,ā and they are shown in a lighter color and with square markers in Figure 2.
@@
Despite the similar model parameters input to both codes, the stellar masses returned by Prospector are larger on average than those from Bagpipes, by approximately an order of magnitude in most cases. This is mainly a consequence of our lack of data longward of rest-frame 500 nm. Bagpipes prefers to fit a very young stellar population while Prospector tends to fit an older stellar population. The two fits diverge beyond rest-frame 1 m, but without longer-wavelength data to constrain the shape of the SED we cannot rule out either possibility. Bagpipes fits the Lyman-alpha break with a damping wing, while in some cases Prospector prefers Lyman-alpha emission and additional dust extinction in the UV (see figures in the Appendix). Both codes return similar relative between sources, though the Prospector values are shifted slightly higher overall (by % on average, due to fitting the rest-UV with more dust). The best-fit star formation rates from both codes agree roughly with the star formation rate estimated from the F405N - F360M flux, i.e. from the [OĀ III] luminosity. For our purposes, the most important parameter is the stellar mass: we discuss the effect of the two codesā mass discrepancy on our results in Section 5.3.
@@ 4.1 HST-only Candidates
LBG1 is not detected in the F405N segmentation map, meaning it is not included in our narrow band-selected sample. Nonetheless, its observed SED shape is consistent with a galaxy (see Figure 3): Bagpipes fits it at , indicating that while LBG1 is not a low- contaminant, it is probably not associated with the quasar.
LBG2 is robustly detected by Spitzer at 3.6 and 4.5 m. Simpson etĀ al. (2014) fit the SED of LBG2, including the Spitzer data, with EAzY, which indicated LBG2 likely lies at rather than . Our JWST data shows a steeply rising SED between F814W and F410M, unlike the rest of our candidates. LBG2 is extremely bright in the redder NIRCam bands (Jy in F405N and F410M, see top right of Figure 3), consistent with its Spitzer detections (though the IRAC1 flux of LBG1 is higher than our F360M flux). Because of this very red SED, LBG2 is selected by our narrow-band excess criterion, but is removed from our final candidate sample because of its large flux and extended shape in the redder NIRCam filters. Bagpipes places LBG2 at .
@@ 5.1 Black Hole Mass
These estimates are widely spread, but can nonetheless help to constrain the Eddington ratio. The theoretical models of Pacucci & Narayan (2024) show that Eddington ratios between 1 and 2.5 lead to for quasars, much higher than observed for J1120+0641. Their models yield lower for high Eddington ratios only if the black hole is viewed pole-on and has significant spin, but these conditions are unlikely for J1120+0641 given the lack of beaming. This implies that J1120+0641 is not accreting significantly above the Eddington limit.
Similarly, the models of Inayoshi etĀ al. (2024) (see their Figure 4) place J1120+0641 below unity Eddington ratio (i.e., ), implying a mildly sub-Eddington accretion rate.
@@ 5.2.1 Far infrared emission and associated SFR
The J1120+0641 host has a substantial far-infrared (far-IR) luminosity, indicating a very high star formation rate. Venemans etĀ al. (2020) use a rest-frame 1900 GHz (158 m) continuum measurement to derive a far-IR luminosity, assuming a modified blackbody model for the far-IR SED, with a temperature of 47K and = 1.6. They multiplied this FIR estimate by 1.41 to correct it to total infrared (TIR) luminosity. However, a modified blackbody may not be the best model at such high redshifts, and the FIR SED of low-redshift, low-metallicity Lyman continuum emitter Haro 11 should provide a more accurate model for the host (see e.g. Lyu etĀ al., 2016; De Rossi etĀ al., 2018). We therefore normalize the SED of Haro 11 to the 158 m measurement and correct to TIR by integrating, finding a correction of a factor of 1.63 (rather than 1.41) and a ULIRG-level TIR luminosity of Lā, implying an obscured SFR of 280 Mā yr-1 (Kennicutt & Evans, 2012; Murphy etĀ al., 2011; Hao etĀ al., 2011).
@@
We seek to confirm this estimate by comparing it to other sources. Marshall etĀ al. (2024) estimate a SFR of 12.5 Mā yr-1 from , but because of artifacts from their PSF subtraction, the region within a radius of 02 is not included in this measurementāwe expect this region to contain most of the central starburst, based on the spatial extent of the far-IR emission. The far-IR emission from the J1120+0641 host is confined mainly to the central 1 - 1.5 kpc (Venemans etĀ al., 2017), so this measurement likely misses most of the star formation. Bosman etĀ al. (2024) measured the broad H line (and longer wavelength recombination lines) in an aperture of radius 03 centered on the nucleus, providing a possibility to fill in the missing area with lines less subject to extinction. In FigureĀ 5, we show the result of refitting their line profile, including two broad components as in their fit but adding an unresolved line. This unresolved component reduces the of the fit, and is best fit when offset to the red by 200 km s-1, consistent with the result in SectionĀ 5.1 given that we are probing different regions in the galaxy. The estimated flux of the narrow H line is W m-2. The error was estimated by repeating the fit with the narrow component constrained to different wavelengths. A similar fit to the Pa profile, keeping the line widths and relative wavelengths (in velocity space) the same but varying the strengths, also reduces the . The line flux is W m-2, with the error determined as for the H line. This is a marginal detection, but is consistent with the expected hydrogen spectrum of a strongly obscured ULIRG.
@@
The ratio of the narrow components of Pa and H implies strong extinction. The extinction-corrected SFR deduced from the H is about 40 Mā yr-1. Combining it with the Marshall etĀ al. (2024) H measurement at larger radii yields a total of 50 Mā yr-1, less than 20% of the FIR SFR. This difference must be attributed to very strong extinction in the star-forming region, supported by the compactness and strength of the far-IR emission. Such behavior is typical of local ULIRGs.
@@
The host galaxy of J1120+0641 therefore appears very similar to lower redshift ULIRGs of similar luminosity. The typical timescale for this level of star forming activity is less than 10 Myr (e.g., Thornley etĀ al., 2000; Bekki etĀ al., 2006; Kawaguchi etĀ al., 2020). Thus the current burst of star formation can be expected to add about Mā to the host galaxy mass. However, these newly forming stars are hidden by the dust extinction from the rest optical and UV (wavelengths probed by HST and JWST observations of the quasar) so we can estimate the mass of the older stars without folding in the effects of these very young ones.

### v3

@@ 1 Introduction
However, the fact remains that such extreme galaxies, with , are not observed in the local Universe aside from a few niche cases where the host galaxy has lost significant mass to tidal stripping. If these quasar host galaxies are to grow enough to land on the observed local relationship between SMBH mass and galaxy mass by , they must be able to form stars and merge with other galaxies in their environment. This fact, combined with predictions of the CDM paradigm and results of cosmological simulations, means when we observe quasars at the end of the Epoch of Reionization, we expect to see them living in dense environments, surrounded by other galaxies (Overzier etĀ al., 2009a; Inayoshi etĀ al., 2020; Volonteri etĀ al., 2021).
For decades, observations have targeted high-redshift quasars in hopes of detecting galaxy overdensities surrounding them. At high redshift (), HST and ground-based optical/NIR telescopes can only access the rest-frame ultraviolet, and must rely on the Lyman break technique (Adams etĀ al., 2015; Husband etĀ al., 2013; Morselli etĀ al., 2014; Ota etĀ al., 2018; Pudoka etĀ al., 2024; Stiavelli etĀ al., 2005; Zheng etĀ al., 2006) or selection of Lyman-alpha emitters (Chanchaiworawit etĀ al., 2019; Harikane etĀ al., 2019; Hennawi etĀ al., 2015; Kashikawa etĀ al., 2007; Overzier etĀ al., 2009b), with relatively few bands longward of the Lyman break, to identify candidate galaxies. Although galaxies were detected around many quasars, it remained difficult to characterize the environments of high-redshift quasars without access to the rest-frame optical at high sensitivity and resolution.
@@ 2.1 JWST NIRCam
We pair each of these NIRCam LW filters with a SW filter. F090W ( = 10522.0 s) straddles the Lyman break at the quasar redshift, and together with the deep archival HST F814W observations (see Section 2.2) will add valuable Lyman-break redshift constraints to exclude lower- and higher-redshift interlopers. Of particular concern is H, which enters the F405N bandpass at : these galaxies may be easily confused for [OĀ III] emitters. We therefore also observe in F150W ( = 5239.5 s) and F200W ( = 5239.5 s), which fall on either side of the Balmer break at and allow us to more easily identify and remove H contaminants.
provide additional photometric redshift constraints to exclude lower- and higher-redshift interlopers from our sample. In particular, F090W straddles the Lyman break and F814W lies blueward of it for galaxies near .
@@ 3.1 Candidate Identification
We visually inspected these 523 candidates to narrow the sample before performing SED fitting. 69 of the 523 sources are spurious: these are a mix of detector artifacts, shredded diffraction spikes of bright point sources in the image, and real galaxies only partially overlapping with the image in one or more bands (see the top panels of Figure 1 for an example). We also remove 112 galaxies obviously at very low redshift based on clearly resolved spiral structure/dust lanes (88 galaxies). Galaxies at are very unlikely to have diameters (defined as twice their effective radii) (Raikov etĀ al., 2025), so we impose a conservative cut at spatial extent greater than (103 galaxies). Finally, we perform a flux cut to remove any with fluxes inconsistent with even the most extreme galaxy (1 Jy in the UV continuum bands F090W, F150W, and/or F200W) not discarded in the visual inspection. This removes another 48 galaxies. For examples of discarded galaxies, see the lower panels of Figure 1. These low-redshift galaxies are sufficiently different from possible contaminants in our study (in SED as well as morphology) that they do not need further study.
@@ 3.2 Spectral Energy Distribution Fitting
For two independent determinations of the exact redshifts of our 294 candidates, we pass them through two SED fitting codes: Prospector and Bagpipes.
@@
We ran Prospector under these conditions three times for the 294 candidates identified in Section 3.1, with different tophat redshift priors: (v1), (v2), and finally (v3; the [OĀ III] line shifts into the F405N bandpass for this redshift range). Galaxies far from the quasar redshift will be worse-fit under the stricter redshift priors, yielding a larger . On the other hand, galaxies near the quasar redshift are likely to have a smaller value in v3 than in v2 or v1, and/or be fit to in the less restrictive v1 or v2 runs.
@@
We identify 32 sources with a smaller with the most restrictive redshift prior (v3) than in v2 or v1. We also find 5 sources that have higher values in v3, but are also fit to in the lower v1 or v2 runs. Our Prospector-curated sample therefore consists of 37 galaxies total.
@@
We also fit the 294 candidates with Bagpipes. We ran Bagpipes with a delayed-tau star formation history and Kroupa (2001) initial mass function, as with Prospector. The Kriek & Conroy (2013) dust law is not supported in Bagpipes, so we use the Salim etĀ al. (2018) law, which is parameterized similarly. We allow the maximum age of the stellar population to vary from zero to the age of the Universe at the fit redshift. The total mass of stars formed was allowed to vary from 106 to 1012 Mā, and the metallicity from 0.1% to 100% solar. Modifying the functional form of the star formation history did not significantly affect the resulting stellar mass (5-10% difference) or quality of the fit. For most sources, modifying the dust extinction law to a CF00 or Calzetti law changed the resulting stellar mass by , though in a few cases the mass changed by up to . As with Prospector, we ran Bagpipes three times with three different redshift ranges: (v1), (v2), and finally (v3). The v3 range is larger for Bagpipes than for Prospector, because Bagpipes imposes a hard limit at the edges of the range, while Prospector can fit galaxies to a redshift outside the prior range.
@@
Bagpipes fits 24 galaxies to a v3 lower than or comparable to v2 and v1; 22 of these candidates overlap with the Prospector sample. Of the 15 galaxies fit to in Prospector but not Bagpipes, nine are only mildly elevated in F405N relative to F360M and F410M, and Bagpipes fits them to based on the position of their Lyman break. Three more are fit to , with F405N and/or F410M on H, and the remaining three are fit to . These galaxies also tend to have larger values in the Prospector fits, and tend to lie at the bright end of the Prospector sample.
We therefore proceed with the sample of 22 galaxies which both Prospector and Bagpipes robustly fit to . Their properties are summarized in Table 1. We note that the redshift range probed by the F405N filter, , is not perfectly centered at the quasar redshift (), which may introduce a sampling effect. We do not believe this makes a significant difference in our final sample, as the vast majority of our final sample have best-fit redshifts near the center of this range.
Of the 272 candidates not robustly localized to , the majority () fall into one of two categories. 118 candidates are fit by both Prospector and Bagpipes as Lyman-break galaxies at high redshift. These galaxies generally display a flat spectrum longward of F090W, and decline sharply in F090W before disappearing altogether in F814W. Another 50 galaxies are fit to by both codes. These galaxies tend to be very red across our observed bands, and are fit with older stellar populations (stellar age generally Gyr) and . These galaxies are bright in F360M, F405N, and F410M, which sit on the peak of the old stellar emission at 1.5-2 m, and the flux declines at shorter wavelengths. This decline in the bluer bands can mimic a Lyman break: a further 18 galaxies are fit to by Bagpipes but with a Lyman break () by Prospector.
@@ 3.3 Final Sample
| ID | RA (ā) | Dec (ā) | Field | aaThe uncertainties on the redshifts are small due to the narrow-band selection method; the presence of the bright F405N and the Lyman break in F090W and F814W robustly places the high-confidence candidates between . The position of the Lyman break is not as clear for the three medium-confidence candidates (see Figure 11), and they are generally not as bright in F405N relative to F360M. Their redshift errors are slightly larger, closer to . | aaThe uncertainties on the redshifts are small due to the narrow-band selection method; the presence of the bright F405N and the Lyman break in F090W and F814W robustly places the high-confidence candidates between . The position of the Lyman break is not as clear for the three medium-confidence candidates (see Figure 11), and they are generally not as bright in F405N relative to F360M. Their redshift errors are slightly larger, closer to . | /Mā) | /Mā) | |
| 300 | 169.99902 | 6.70370 | Q | 7.11 | 7.10 | |||
| 335 | 169.99734 | 6.69700 | Q | 7.10 | 7.11 | |||
| 337 | 169.99750 | 6.69704 | Q | 7.11 | 7.11 | |||
| 414 | 169.99737 | 6.68958 | Q | 7.10 | 7.12 | |||
| 427 | 169.99444 | 6.68186 | Q | 7.11 | 7.11 | |||
| 534 | 170.00581 | 6.69992 | Q | 7.11 | 7.11 | |||
| 624 | 170.00959 | 6.70138 | Q | 7.11 | 7.11 | |||
| 666 | 170.00589 | 6.68920 | Q | 7.11 | 7.11 | |||
| 703 | 170.00662 | 6.68699 | Q | 7.11 | 7.11 | |||
| 709 | 170.00678 | 6.68705 | Q | 7.07 | 7.10 | |||
| 789 | 170.01207 | 6.69086 | Q | 7.14 | 7.11 | |||
| 876 | 170.01181 | 6.68086 | Q | 7.10 | 7.11 | |||
| 941 | 170.01047 | 6.67159 | Q | 7.13 | 7.11 | |||
| 181 | 169.97607 | 6.66299 | A | 7.11 | 7.08 | |||
| 358 | 169.97943 | 6.65433 | A | 7.11 | 7.08 | |||
| 366 | 169.97875 | 6.65236 | A | 7.11 | 7.08 | |||
| 456 | 169.97744 | 6.64111 | A | 7.12 | 7.08 | |||
| 634 | 169.97744 | 6.62748 | A | 7.09 | 7.08 | |||
| 693 | 169.98564 | 6.64099 | A | 7.11 | 7.08 | |||
| 160* | 169.96941 | 6.64997 | A | 7.23 | 7.31 | |||
| 856* | 169.99940 | 6.65472 | A | 7.01 | 7.41 | |||
| 911* | 169.98969 | 6.62826 | A | 7.20 | 7.17 |
The positions of the 22 candidates are shown, along with the position of the quasar, in Figure 2, and their thumbnails and SEDs are plotted in the Appendix. Thirteen candidates lie in Module B, in the same module as the quasar; we hereafter refer to Module B as the āquasar field.ā The remaining nine candidates lie in Module A, hereafter the āadjacent fieldā (though note that an overdensity around J1120+0641 is likely to span both fields). The vast majority of these candidates are fit to a redshift between 7.07 and 7.14 by both Prospector and Bagpipes, but three of the candidates in the adjacent field are fit farther from the quasar redshift by both codes. We consider these three candidates āmedium confidence,ā and they are shown in a lighter color and with square markers in Figure 2.
@@
Despite the similar model parameters input to both codes, the stellar masses returned by Prospector are larger on average than those from Bagpipes, by approximately an order of magnitude in most cases. This is mainly a consequence of our lack of data longward of rest-frame 500 nm. Bagpipes prefers to fit a very young stellar population while Prospector tends to fit an older stellar population. The two fits diverge beyond rest-frame 1 m, but without longer-wavelength data to constrain the shape of the SED we cannot rule out either possibility. Bagpipes fits the Lyman-alpha break with a damping wing, while in some cases Prospector prefers Lyman-alpha emission and additional dust extinction in the UV (see figures in the Appendix). Both codes return similar relative between sources, though the Prospector values are shifted slightly higher overall (by % on average, due to fitting the rest-UV with more dust). The best-fit star formation rates (SFRs) from both codes agree roughly with the SFR estimated from the F405N - F360M flux, i.e. from the [OĀ III] luminosity. For our purposes, the most important parameter is the stellar mass: we discuss the effect of the two codesā mass discrepancy on our results in Section 5.3.
@@ 4.1 HST-only Candidates
LBG1 is not detected in the F405N segmentation map, meaning it is not included in our narrow band-selected sample and its redshift determination is less certain. Its observed SED shape is consistent with a galaxy (see Figure 3), and Bagpipes fits it at , primarily from the position of the Lyman break; given the width of the F090W bandpass, the uncertainty on this redshift is of order . This is in contrast to the high-confidence overdensity candidates, where the F405N excess combined with the Lyman break allows them to be fit to with much smaller redshift uncertainties. LBG1 is certainly not a low- contaminant, but without an F405N excess we cannot confirm whether it is directly associated with the quasar and simply [OĀ III]-faint, or an unrelated galaxy at slightly lower or higher redshift than the quasar. Because of its uncertain redshift, we do not include LBG1 in the candidate sample going forward.
LBG2 is robustly detected by Spitzer at 3.6 and 4.5 m. Simpson etĀ al. (2014) fit the SED of LBG2, including the Spitzer data, with EAzY, which indicated LBG2 likely lies at rather than . Our JWST data shows a steeply rising SED between F814W and F410M, unlike the rest of our candidates. LBG2 is extremely bright in the redder NIRCam bands (Jy in F405N and F410M, see top right of Figure 3), consistent with its Spitzer detections (though the IRAC1 flux of LBG1 is higher than our F360M flux). Because of this very red SED, LBG2 is selected by our narrow-band excess criterion, but is removed from our final candidate sample at the visual inspection stage because of its large flux and extended shape in the redder NIRCam filters. Bagpipes places LBG2 at .
@@ 5.1 Black Hole Mass
These estimates are widely spread, but nonetheless indicate that the luminosity is not strongly super-Eddington. The theoretical models of Pacucci & Narayan (2024) show that Eddington ratios between 1 and 2.5 lead to for quasars, much higher than indicated for J1120+0641. Their models yield lower for high Eddington ratios only if the black hole is viewed pole-on and has significant spin, but these conditions are unlikely for J1120+0641 given the lack of beaming. This implies that J1120+0641 is not accreting significantly above the Eddington limit.
Similarly, the models of Inayoshi etĀ al. (2024) (see their Figure 4) place J1120+0641 below unity Eddington ratio (i.e., ). Allowing for errors, a highly super-Eddington ratio is not plausible, although a mildly super-Eddington one is possible.
@@ 5.2.1 Far infrared emission and associated SFR
The J1120+0641 host has a substantial far-infrared (far-IR) luminosity, indicating a very high SFR. Venemans etĀ al. (2020) use a rest-frame 1900 GHz (158 m) continuum measurement to derive a far-IR luminosity, assuming a modified blackbody model for the far-IR SED, with a temperature of 47K and = 1.6. They multiplied this FIR estimate by 1.41 to correct it to total infrared (TIR) luminosity. However, a modified blackbody may not be the best model at such high redshifts, and the FIR SED of low-redshift, low-metallicity Lyman continuum emitter Haro 11 should provide a more accurate model for the host (see e.g. Lyu etĀ al., 2016; De Rossi etĀ al., 2018). We therefore normalize the SED of Haro 11 to the 158 m measurement and correct to TIR by integrating, finding a correction of a factor of 1.63 (rather than 1.41) and a ULIRG-level TIR luminosity of Lā, implying an obscured SFR of 280 Mā yr-1 (Kennicutt & Evans, 2012; Murphy etĀ al., 2011; Hao etĀ al., 2011).
@@
In general, ULIRGs absorb most of the energy in emission lines due to young stars, with only a fraction that escapes, meaning SFR estimates from optical and near-IR emission lines underestimate the total star formation. We can check for this behavior by comparing J1120+0641ās SFR from the infrared luminosity with what would be deduced from standard optical emission line estimates. The much lower SFR from H (12.5 Mā yr-1, Marshall etĀ al., 2024) seems to support this point, but because of artifacts from their PSF subtraction, the region within a radius of 02 is not included in this measurement. We expect this region to contain most of the central starburst, based on the 1 - 1.5 kpc spatial extent of the far-IR emission (Venemans etĀ al., 2017). Bosman etĀ al. (2024) measured H in an aperture of radius 03 centered on the nucleus, providing a possibility to fill in the missing area. In FigureĀ 5, we show the result of refitting their line profile, including two broad components as in their fit but adding an unresolved line. This unresolved component reduces the of the fit, based on an error estimated by repeating the fit multiple times with the narrow component set to different wavelengths. This fit is offset to the red by 200 km s-1, consistent with the result in SectionĀ 5.1 given that we are probing different regions in the galaxy. The estimated flux of the narrow H line is W m-2. A similar fit to the Pa profile, keeping the line widths and relative wavelengths (in velocity space) shows a weak detection (). The fact that neither H nor Pa show a strong narrow line supports the host galaxyās ULIRG behavior, indicating strong obscuration that suppresses optical/near-IR star formation signatures.
@@
The SFR deduced from the H is about 40 Mā yr-1. Combining it with the Marshall etĀ al. (2024) H measurement at larger radii yields a total of Mā yr-1, less than 20% of the FIR SFR. This difference must be attributed to very strong extinction in the star-forming region, supported by the compactness and strength of the far-IR emission. Such behavior is typical of high redshift ULIRGs (Swinbank etĀ al., 2004; Takata etĀ al., 2006).
@@
The host galaxy of J1120+0641 therefore appears very similar to lower redshift ULIRGs of similar luminosity. The typical timescale for this level of star forming activity is less than 10 Myr (e.g., Thornley etĀ al., 2000; Bekki etĀ al., 2006; Kawaguchi etĀ al., 2020). Thus the current burst of star formation can be expected to add about Mā to the host galaxy mass. These newly forming stars are hidden by the dust extinction from the rest optical and UV. We can therefore use existing HST and JWST observations of the quasar to characterize the older stellar population that dominates the mass.